OpenStudy (kakarot):

Medal! Solve for r 9(r – 82) – 5 = 58

OpenStudy (kakarot):

Its kinda hard

OpenStudy (dayakar):

multiply inside the bracket terms with 9 remaining terms are same

OpenStudy (kakarot):

r=89?

OpenStudy (dayakar):

no

OpenStudy (kakarot):

what? we are solving for r. what does r equal?

OpenStudy (dayakar):

r is called a variable means unknown value

OpenStudy (kakarot):

I know but the answer to the equation has to be 89

OpenStudy (kakarot):

r=89

OpenStudy (dayakar):

how can u get it?

OpenStudy (kakarot):

expand, add 743 to both sides, divide both sides by 9. r=89

OpenStudy (dayakar):

9(r-82)-5 =58 9(r-82) =58+5 9(r-82)= 63

OpenStudy (kakarot):

your math is wrong

OpenStudy (kakarot):

@Kainui

OpenStudy (dayakar):

your answer is right , u need to write steps

OpenStudy (kakarot):

Oh. I dont need to right the steps for this specific question because it just want the anwser for r

OpenStudy (parker.goodbar):

i suggest you write the steps so you can refer back to it for later questions

OpenStudy (dayakar):

r-82 = 63/9 r-82 = 7 r=7+82 r=89

OpenStudy (kakarot):

I did write the steps. Im working in IXL so if I get it wrong it will explain it to me and send me back like 5 questions and make me redo it

OpenStudy (kakarot):

with different questions of course. and it doesnt effect my grade

OpenStudy (parker.goodbar):

oh ok cool